OpenStudy (anonymous):

What is the solution to the equation log4 x – log4 2 = 3 ?

OpenStudy (anonymous):

help anyone

OpenStudy (anonymous):

Use The Property LogX-LogY=Log(X/y)

OpenStudy (anonymous):

Any Hint ?

OpenStudy (anonymous):

show me an example

hartnn (hartnn):

try to use this formula: \[\log_{4}A-\log_{4}B=\log_{4}\frac{A}{B} \]

OpenStudy (anonymous):

x/2?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

can u start it and i finish

hartnn (hartnn):

ok,so using that formula,tell me whats \[\log_{4}x-\log_{4}2 =? \]

OpenStudy (anonymous):

log4 x/2?

hartnn (hartnn):

yup :) now lets go ahead...

OpenStudy (anonymous):

so now what do u do

hartnn (hartnn):

if \[\log_{x}y=z \] then \[x^{z}=y\]

hartnn (hartnn):

here u have \[\log_{4}(x/2)=3 \] so what do u get from this?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

do u multpy something or what

OpenStudy (anonymous):

eyad help please

OpenStudy (anonymous):

Change it To exponential .... 4^3=x/2 64=x/2 =====> multiply by 2 x=128

OpenStudy (anonymous):

Got It ?

OpenStudy (anonymous):

oh ok thnx so much now i understand

OpenStudy (anonymous):

COoOL ,Yw ..